### v4

@@ 2 Proposed Bayesian model for variable selection in functional regression
2.1 Full conditional distributions
$f(\boldsymbol{b}|\boldsymbol{\theta},\boldsymbol{\mu},\sigma^{2},\boldsymbol{\tau^{2}},\boldsymbol{Z},\boldsymbol{y})$
$\propto\pi(\sigma^{2})\pi(\boldsymbol{\tau^{2}})\pi(\boldsymbol{\mu})f(\boldsymbol{\theta}|\boldsymbol{\mu})f(\boldsymbol{b}|\sigma^{2},\boldsymbol{\tau^{2}})p(\boldsymbol{Z}|\boldsymbol{\theta})f(\boldsymbol{y}|\boldsymbol{b},\boldsymbol{Z},\sigma^{2})$
$\propto\left[\prod_{k=1}^{K}\prod_{l=1}^{p}f(b_{kl}|\sigma^{2},\tau_{kl}^{2})\right]\left[\prod_{i=1}^{m}\prod_{j=1}^{n_{i}}f(y_{ij}|\boldsymbol{b},\boldsymbol{Z},\sigma^{2})\right]$
$\propto\exp\left\{-\frac{\sum_{k=1}^{K}\sum_{l=1}^{p}\frac{b_{kl}^{2}}{\tau_{kl}^{2}}}{2\sigma^{2}}\right\}\exp\left\{-\frac{\sum_{i=1}^{m}\sum_{j=1}^{n_{i}}(\tilde{y}_{ij}-g_{i}(t_{ij}))^{2}}{2\sigma^{2}}\right\}\text{,}$
in which $\tilde{y}_{ij}=y_{ij}-\beta_{0}(t_{ij})$.
Remembering that $g_{i}(t_{ij})=\sum_{l=1}^{p}\sum_{k=1}^{K}x_{li}Z_{l}b_{kl}B_{k}(t_{ij})$, considering
$\boldsymbol{O_{lij}}=(x_{li}Z_{l}B_{1}(t_{ij}),\;x_{li}Z_{l}B_{2}(t_{ij}),\dots,\;x_{li}Z_{l}B_{K}(t_{ij}))^{{}^{\prime}}\text{,}$
a dimension vector $(K\times 1)$, and calculating $\boldsymbol{O_{.ij}}=\begin{pmatrix}\boldsymbol{O_{1ij}}^{{}^{\prime}},&\boldsymbol{O_{2ij}}^{{}^{\prime}},&\dots,&\boldsymbol{O_{pij}^{{}^{\prime}}}\end{pmatrix}_{(Kp\times 1)}^{{}^{\prime}}$, so we have $g_{i}(t_{ij})=\boldsymbol{b}^{{}^{\prime}}\boldsymbol{O_{.ij}}$.
As such, $\boldsymbol{\eta^{2}}=(\boldsymbol{\eta^{2}_{.1}}^{{}^{\prime}},\;\boldsymbol{\eta^{2}_{.2}}^{{}^{\prime}},\dots,\;\boldsymbol{\eta^{2}_{.p}}^{{}^{\prime}})^{{}^{\prime}}$, in which $\boldsymbol{\eta^{2}_{.l}}=(\frac{1}{\tau^{2}_{1l}},\frac{1}{\tau^{2}_{2l}},\dots,\frac{1}{\tau^{2}_{Kl}})^{{}^{\prime}}$, so we have
$f(\boldsymbol{b}|\boldsymbol{\theta},\boldsymbol{\mu},\sigma^{2},\tau^{2},\boldsymbol{Z},\boldsymbol{y})$
$\propto\exp\left\{-\frac{\boldsymbol{b}^{{}^{\prime}}\operatorname*{diag}(\boldsymbol{\eta^{2}})\boldsymbol{b}+\sum_{i=1}^{m}\sum_{j=1}^{n_{i}}(\tilde{y}_{ij}-\boldsymbol{b}^{{}^{\prime}}\boldsymbol{O_{.ij}})^{{}^{\prime}}(\tilde{y}_{ij}-\boldsymbol{b}^{{}^{\prime}}\boldsymbol{O_{.ij}})}{2\sigma^{2}}\right\}$
$\propto\exp\left\{-\frac{\boldsymbol{b}^{{}^{\prime}}\operatorname*{diag}(\boldsymbol{\eta^{2}})\boldsymbol{b}+\sum_{i=1}^{m}\sum_{j=1}^{n_{i}}\left[\tilde{y}_{ij}^{2}-2\tilde{y}_{ij}\boldsymbol{b}^{{}^{\prime}}\boldsymbol{O_{.ij}}+\boldsymbol{b}^{{}^{\prime}}\boldsymbol{O_{.ij}}\boldsymbol{O_{.ij}}^{{}^{\prime}}\boldsymbol{b}\right]}{2\sigma^{2}}\right\}$
$\propto\exp\left\{-\frac{\boldsymbol{b}^{{}^{\prime}}\left[\operatorname*{diag}(\boldsymbol{\eta^{2}})+\sum_{i=1}^{m}\sum_{j=1}^{n_{i}}\boldsymbol{O_{.ij}}\boldsymbol{O_{.ij}}^{{}^{\prime}}\right]\boldsymbol{b}-2\boldsymbol{b}^{{}^{\prime}}\sum_{i=1}^{m}\sum_{j=1}^{n_{i}}\boldsymbol{O_{.ij}}\tilde{y}_{ij}}{2\sigma^{2}}\right\}\text{.}$
By the Cholesky decomposition, let $\boldsymbol{Q}=\left[\operatorname*{diag}(\boldsymbol{\eta^{2}})+\sum_{i=1}^{m}\sum_{j=1}^{n_{i}}\boldsymbol{O_{.ij}}\boldsymbol{O_{.ij}}^{{}^{\prime}}\right]=\boldsymbol{L_{i}}\boldsymbol{L_{i}}^{{}^{\prime}}$, a matrix of dimension $(Kp\times Kp)$, then
$f(\boldsymbol{b}|\boldsymbol{\theta},\boldsymbol{\mu},\sigma^{2},\tau^{2},\boldsymbol{Z},\boldsymbol{y})$
$\propto\exp\left\{-\frac{\boldsymbol{b}^{{}^{\prime}}\boldsymbol{L_{i}}\boldsymbol{L_{i}}^{{}^{\prime}}\boldsymbol{b}-2\boldsymbol{b}^{{}^{\prime}}\sum_{i=1}^{m}\sum_{j=1}^{n_{i}}\boldsymbol{O_{.ij}}\tilde{y}_{ij}}{2\sigma^{2}}\right\}$
$\propto\exp\left\{-\frac{(\boldsymbol{L_{i}}^{{}^{\prime}}\boldsymbol{b})^{{}^{\prime}}\boldsymbol{L_{i}}^{{}^{\prime}}\boldsymbol{b}-2\boldsymbol{b}^{{}^{\prime}}\boldsymbol{L_{i}}\boldsymbol{L_{i}}^{-1}\sum_{i=1}^{m}\sum_{j=1}^{n_{i}}\boldsymbol{O_{.ij}}\tilde{y}_{ij}}{2\sigma^{2}}\right\}$
$\propto\exp\left\{-\frac{(\boldsymbol{L_{i}}^{{}^{\prime}}\boldsymbol{b})^{{}^{\prime}}\boldsymbol{L_{i}}^{{}^{\prime}}\boldsymbol{b}-2(\boldsymbol{L_{i}}^{{}^{\prime}}\boldsymbol{b})^{{}^{\prime}}\boldsymbol{L_{i}}^{-1}\sum_{i=1}^{m}\sum_{j=1}^{n_{i}}\boldsymbol{O_{.ij}}\tilde{y}_{ij}}{2\sigma^{2}}\right\}$
$\propto\exp\left\{-\frac{(\boldsymbol{L_{i}}^{{}^{\prime}}\boldsymbol{b}-\boldsymbol{L_{i}}^{-1}\sum_{i=1}^{m}\sum_{j=1}^{n_{i}}\boldsymbol{O_{.ij}}\tilde{y}_{ij})^{{}^{\prime}}(\boldsymbol{L_{i}}^{{}^{\prime}}\boldsymbol{b}-\boldsymbol{L_{i}}^{-1}\sum_{i=1}^{m}\sum_{j=1}^{n_{i}}\boldsymbol{O_{.ij}}\tilde{y}_{ij})}{2\sigma^{2}}\right\}$
$\propto\exp\left\{-\frac{(\boldsymbol{L_{i}}^{{}^{\prime}}\boldsymbol{b}-\boldsymbol{L_{i}}^{-1}\sum_{i=1}^{m}\sum_{j=1}^{n_{i}}\boldsymbol{O_{.ij}}\tilde{y}_{ij})^{{}^{\prime}}\boldsymbol{L_{i}}^{-1}\boldsymbol{L_{i}}\boldsymbol{L_{i}}^{{}^{\prime}}(\boldsymbol{L_{i}}^{{}^{\prime}})^{-1}(\boldsymbol{L_{i}}^{{}^{\prime}}\boldsymbol{b}-\boldsymbol{L_{i}}^{-1}\sum_{i=1}^{m}\sum_{j=1}^{n_{i}}\boldsymbol{O_{.ij}}\tilde{y}_{ij})}{2\sigma^{2}}\right\}$
$\propto\exp\left\{-\frac{(\boldsymbol{b}-(\boldsymbol{L_{i}}\boldsymbol{L_{i}}^{{}^{\prime}})^{-1}\sum_{i=1}^{m}\sum_{j=1}^{n_{i}}\boldsymbol{O_{.ij}}\tilde{y}_{ij})^{{}^{\prime}}\boldsymbol{L_{i}}\boldsymbol{L_{i}}^{{}^{\prime}}(\boldsymbol{b}-(\boldsymbol{L_{i}}\boldsymbol{L_{i}}^{{}^{\prime}})^{-1}\sum_{i=1}^{m}\sum_{j=1}^{n_{i}}\boldsymbol{O_{.ij}}\tilde{y}_{ij})}{2\sigma^{2}}\right\}\text{.}$
As a result,
$f(\boldsymbol{b}|\boldsymbol{\theta},\boldsymbol{\mu},\sigma^{2},\boldsymbol{\tau^{2}},\boldsymbol{Z},\boldsymbol{y})=f(\boldsymbol{b}|{\sigma^{2}},\boldsymbol{\tau^{2}},\boldsymbol{Z},\boldsymbol{y})$
$\propto\exp\left\{-\frac{(\boldsymbol{b}-\boldsymbol{Q}^{-1}\sum_{i=1}^{m}\sum_{j=1}^{n_{i}}\boldsymbol{O_{.ij}}\tilde{y}_{ij})^{{}^{\prime}}(\boldsymbol{Q}^{-1})^{-1}(\boldsymbol{b}-\boldsymbol{Q}^{-1}\sum_{i=1}^{m}\sum_{j=1}^{n_{i}}\boldsymbol{O_{.ij}}\tilde{y}_{ij})}{2\sigma^{2}}\right\}\text{.}$ (6)
So, $f(\boldsymbol{b}|\boldsymbol{\theta},\boldsymbol{\mu},\sigma^{2},\boldsymbol{\tau^{2}},\boldsymbol{Z},\boldsymbol{y})$ is precisely a Multivariate Normal distribution with mean $\boldsymbol{Q}^{-1}\sum_{i=1}^{m}\sum_{j=1}^{n_{i}}\boldsymbol{O_{.ij}}\tilde{y}_{ij}$ and variance $\sigma^{2}\boldsymbol{Q}^{-1}$.
The mean of this Normal distribution can still be described in a simplified way. When considering
so we have
Similarly, we obtain the full conditionals of the $\theta_{l}$’s, taking
$f(\theta_{l}|\boldsymbol{b},\boldsymbol{\theta_{-[l]}},\boldsymbol{\mu},\sigma^{2},\boldsymbol{\tau^{2}},\boldsymbol{Z},\boldsymbol{y})$
$\propto\pi(\sigma^{2})\pi(\boldsymbol{\tau^{2}})\pi(\boldsymbol{\mu})f(\theta_{l}|\mu_{l})f(\boldsymbol{\theta_{-[l]}}|\boldsymbol{\mu_{-[l]}})f(\boldsymbol{b}|\sigma^{2},\boldsymbol{\tau^{2}})p(\boldsymbol{Z}|\boldsymbol{\theta})p(\boldsymbol{y}|\boldsymbol{b},\boldsymbol{Z},\sigma^{2})$
$\propto f(\theta_{l}|\mu_{l})p(Z_{l}|\theta_{l})\propto(\theta_{l})^{\mu_{l}-1}(1-\theta_{l})^{(1-\mu_{l})-1}(\theta_{l})^{Z_{l}}(1-\theta_{l})^{1-Z_{l}}\text{.}$
Thus,
$\propto(\theta_{l})^{\mu_{l}+Z_{l}-1}(1-\theta_{l})^{(1-\mu_{l})-Z_{l}+1-1}\text{.}$ (7)
Which shows that $f(\theta_{l}|\boldsymbol{b},\boldsymbol{\theta_{-[l]}},\boldsymbol{\mu},\sigma^{2},\boldsymbol{\tau^{2}},\boldsymbol{Z},\boldsymbol{y})$ is a beta with the first parameter being $\mu_{l}+Z_{l}$ and the second being $2-Z_{l}-\mu_{l}$.
$\scriptsize=\frac{\pi(\sigma^{2})\pi(\boldsymbol{\tau^{2}})\pi(\boldsymbol{\mu})f(\boldsymbol{\theta}|\boldsymbol{\mu})f(\boldsymbol{b}|\sigma^{2},\boldsymbol{\tau^{2}})\mathrm{P}(Z_{l}=1|\boldsymbol{\theta})\mathrm{P}(\boldsymbol{Z_{-[l]}}=z_{-[l]}|\boldsymbol{\theta})f(\boldsymbol{y}|\boldsymbol{b},Z_{l}=1,\boldsymbol{Z_{-[l]}}=z_{-[l]},\sigma^{2})}{\sum_{z=0}^{1}\pi(\sigma^{2})\pi(\boldsymbol{\tau^{2}})\pi(\boldsymbol{\mu})f(\boldsymbol{\theta}|\boldsymbol{\mu})f(\boldsymbol{b}|\sigma^{2},\boldsymbol{\tau^{2}})\mathrm{P}(Z_{l}=z|\boldsymbol{\theta})\mathrm{P}(\boldsymbol{Z_{-[l]}}=z_{-[l]}|\boldsymbol{\theta})f(\boldsymbol{y}|\boldsymbol{b},Z_{l}=z,\boldsymbol{Z_{-[l]}}=z_{-[l]},\sigma^{2})}$
$\normalsize=\frac{\mathrm{P}(Z_{l}=1|\boldsymbol{\theta})\mathrm{P}(\boldsymbol{Z_{-[l]}}=z_{-[l]}|\boldsymbol{\theta})f(\boldsymbol{y}|\boldsymbol{b},Z_{l}=1,\boldsymbol{Z_{-[l]}}=z_{-[l]},\sigma^{2})}{\sum_{z=0}^{1}\mathrm{P}(Z_{l}=z|\boldsymbol{\theta})\mathrm{P}(\boldsymbol{Z_{-[l]}}=z_{-[l]}|\boldsymbol{\theta})f(\boldsymbol{y}|\boldsymbol{b},Z_{l}=z,\boldsymbol{Z_{-[l]}}=z_{-[l]},\sigma^{2})}\text{.}$
Simplifying this, we have
$=\frac{\theta_{l}f(\boldsymbol{y}|\boldsymbol{b},Z_{l}=1,\boldsymbol{Z_{-[l]}}=z_{-[l]},\sigma^{2})}{(1-\theta_{l})f(\boldsymbol{y}|\boldsymbol{b},Z_{l}=0,\boldsymbol{Z_{-[l]}}=z_{-[l]},\sigma^{2})+\theta_{l}f(\boldsymbol{y}|\boldsymbol{b},Z_{l}=1,\boldsymbol{Z_{-[l]}}=z_{-[l]},\sigma^{2})}$
Let $\boldsymbol{B_{ij}}=(B_{1}(t_{ij}),\;B_{2}(t_{ij}),\dots,\;B_{K}(t_{ij}))^{{}^{\prime}}$ and $g_{i,0}(t_{ij})=\sum_{q\neq l}x_{qi}Z_{q}\boldsymbol{b_{.q}}^{{}^{\prime}}\boldsymbol{B_{ij}}$ and $g_{i,1}(t_{ij})=x_{li}\boldsymbol{b_{.l}}^{{}^{\prime}}\boldsymbol{B_{ij}}+\sum_{q\neq l}x_{qi}Z_{q}\boldsymbol{b_{.q}}^{{}^{\prime}}\boldsymbol{B_{ij}}$, so we have
$\frac{f(\boldsymbol{y}|\boldsymbol{b},Z_{l}=0,\boldsymbol{Z_{-[l]}}=z_{-[l]},\sigma^{2})}{f(\boldsymbol{y}|\boldsymbol{b},Z_{l}=1,\boldsymbol{Z_{-[l]}}=z_{-[l]},\sigma^{2})}$
$=\exp\Bigg{\{}\frac{1}{2\sigma^{2}}\Bigg{[}\sum_{j=1}^{n_{i}}(y_{ij}-\beta_{0}(t_{ij})-g_{i,1}(t_{ij}))^{2}+\sum_{s\neq i}\sum_{j=1}^{n_{s}}(y_{sj}-\beta_{0}(t_{sj})-g_{i}(t_{sj}))^{2}$
$-\sum_{j=1}^{n_{i}}(y_{ij}-\beta_{0}(t_{ij})-g_{i,0}(t_{ij}))^{2}-\sum_{s\neq i}\sum_{j=1}^{n_{s}}(y_{sj}-\beta_{0}(t_{sj})-g_{i}(t_{sj}))^{2}\Bigg{]}\Bigg{\}}$
Then,
$\mathrm{P}(Z_{l}=1|\boldsymbol{b},\boldsymbol{\theta},\boldsymbol{\mu},\sigma^{2},\boldsymbol{\tau^{2}},\boldsymbol{Z_{-[l]}},\boldsymbol{y})=\mathrm{P}(Z_{l}=1|\boldsymbol{b},\theta_{l},{\sigma^{2}},\boldsymbol{Z_{-[l]}},\boldsymbol{y})$
$=\frac{\theta_{l}}{(1-\theta_{l})\exp\Bigg{\{}\frac{1}{2\sigma^{2}}\Bigg{[}\sum_{j=1}^{n_{i}}(\tilde{y}_{ij}-g_{i,1}(t_{ij}))^{2}-\sum_{j=1}^{n_{i}}(\tilde{y}_{ij}-g_{i,0}(t_{ij}))^{2}\Bigg{]}\Bigg{\}}+\theta_{l}}\text{.}$ (8)
The construction of the full conditional distribution of $\sigma^{2}$ is already more straightforward. Taking into account that
$\propto\pi(\sigma^{2})f(\boldsymbol{b}|\sigma^{2},\boldsymbol{\tau^{2}})f(\boldsymbol{y}|\boldsymbol{b},\boldsymbol{Z},\sigma^{2})\propto\pi(\sigma^{2})\left[\prod_{k=1}^{K}\prod_{l=1}^{p}f(b_{kl}|\sigma^{2},\tau_{kl}^{2})\right]\left[\prod_{i=1}^{m}\prod_{j=1}^{n_{i}}f(y_{ij}|\boldsymbol{b},\boldsymbol{Z},\sigma^{2})\right]$
×$\times\left[\prod_{i=1}^{m}\prod_{j=1}^{n_{i}}\left(\frac{1}{\sigma^{2}}\right)^{\frac{1}{2}}\exp\left\{-\frac{(y_{ij}-\beta_{0}(t_{ij})-g_{i}(t_{ij}))^{2}}{2\sigma^{2}}\right\}\right]\text{,}$
We have
$\propto\left(\frac{1}{\sigma^{2}}\right)^{\frac{\sum_{i=1}^{m}n_{i}}{2}+\frac{pK}{2}+\delta_{1}+1}\exp\left\{-\frac{\sum_{i=1}^{m}\sum_{j=1}^{n_{i}}(\tilde{y}_{ij}-g_{i}(t_{ij}))^{2}+\sum_{k=1}^{K}\sum_{l=1}^{p}\frac{b_{kl}^{2}}{\tau_{kl}^{2}}+2\delta_{2}}{2\sigma^{2}}\right\}\text{.}$ (9)
Thus, the full conditional distribution of $\sigma^{2}$ is an inverse gamma, with $\frac{\sum_{i=1}^{m}n_{i}}{2}+\frac{pK}{2}+\delta_{1}$ as the shape parameter and $\frac{\sum_{i=1}^{m}\sum_{j=1}^{n_{i}}(\tilde{y}_{ij}-g_{i}(t_{ij}))^{2}+\sum_{k=1}^{K}\sum_{l=1}^{p}\frac{b_{kl}^{2}}{\tau_{kl}^{2}}+2\delta_{2}}{2}$ as the rate parameter.
Finally, it remains to compute the full conditional for the $\tau_{kl}^{2}$’s. It is known that
$\propto\pi(\tau_{kl}^{2})f(b_{kl}|\sigma^{2},\tau_{kl}^{2})\propto\frac{\lambda^{2}}{2}\exp\left\{-\frac{\lambda^{2}}{2}\tau_{kl}^{2}\right\}\left(\frac{1}{\sigma^{2}\tau_{kl}^{2}}\right)^{\frac{1}{2}}\exp\left\{-\frac{b_{kl}^{2}}{2\sigma^{2}\tau_{kl}^{2}}\right\}\text{.}$
Therefore,
$\propto\left(\frac{1}{\tau_{kl}^{2}}\right)^{\frac{1}{2}}\exp\left\{-\left(\frac{\frac{b_{kl}^{2}}{\sigma^{2}}}{2\tau_{kl}^{2}}+\frac{\lambda^{2}}{2}\tau_{kl}^{2}\right)\right\}\text{.}$ (10)
With this structure, it is possible to prove that $\eta_{kl}^{2}=\frac{1}{\tau_{kl}^{2}}$, when conditioned to the other terms of the hierarchical formulation it has normal-inverse distribution, with parameterization in $\lambda^{{}^{\prime}}$ and $\mu^{{}^{\prime}}$, so that
To verify this, we need to rewrite the Expression (10) in terms of $\eta_{kl}^{2}$, including the Jacobian ($-\frac{1}{(\eta_{kl}^{2})^{2}}$). So, the full conditional of $\eta_{kl}^{2}$ is proportional to
$(\eta_{kl}^{2})^{-\frac{3}{2}}\exp\left\{-\frac{1}{2}\left(\frac{b_{kl}^{2}}{\sigma^{2}}\eta_{kl}^{2}+\frac{\lambda^{2}}{\eta_{kl}^{2}}\right)\right\}$
$\propto(\eta_{kl}^{2})^{-\frac{3}{2}}\exp\left\{-\frac{b_{kl}^{2}\left[(\eta_{kl}^{2})^{2}+\frac{\lambda^{2}\sigma^{2}}{b_{kl}^{2}}\right]}{2\sigma^{2}\eta_{kl}^{2}}\right\}$
As a result,
$\propto(\eta_{kl}^{2})^{-\frac{3}{2}}\exp\left\{-\frac{b_{kl}^{2}\left(\eta_{kl}^{2}-\sqrt{\frac{\lambda^{2}\sigma^{2}}{b_{kl}^{2}}}\right)^{2}}{2\sigma^{2}\eta_{kl}^{2}}\right\}\text{.}$ (11)
Then, $\eta_{kl}^{2}$ has a normal-inverse distribution with $\mu^{{}^{\prime}}=\sqrt{\frac{\lambda^{2}\sigma^{2}}{b_{kl}^{2}}}$ and $\lambda^{{}^{\prime}}=\lambda^{2}$.
Thus, to simulate $\tau_{kl}^{2}$ in the Gibbs sampler, it is enough to first simulate $\eta_{kl}^{2}$ to then obtain $\tau_{kl}^{2}$ through the transformation $\tau_{kl}^{2}=\frac{1}{\eta_{kl}^{2}}$.
An additional account must be performed when considering $\mu_{l}$ as a parameter:
Thus,
$\propto\pi(\mu_{l})f(\theta_{l}|\mu_{l})\propto I_{(0,\psi)}(\mu_{l})\theta_{l}^{\mu_{l}}(1-\theta_{l})^{1-\mu_{l}}\text{.}$ (12)
Then, $f(\mu_{l}|\boldsymbol{b},\boldsymbol{\theta},\boldsymbol{\mu_{-[l]}},\sigma^{2},\boldsymbol{\tau^{2}},\boldsymbol{Z},\boldsymbol{y})$ is a continuous bernoulli distribution with parameter $\theta_{l}$ truncated on the interval $(0,\psi)$.
@@ 2.2 Gibbs sampler
The panel 2.2 presents the standard procedure for implementing the Gibbs sampler in a clear and summarized way.
..1. Define the hyperparameters; ..2. Assign initial chain values to parameters; ..3. For $\text{c}=2,\dots,\text{Nint}$: ..3.1 Sample ${\sigma^{2}}^{(c)}\sim f(\sigma^{2}|\boldsymbol{b}^{(c-1)},\boldsymbol{\tau^{2}}^{(c-1)},\boldsymbol{Z}^{(c-1)},\boldsymbol{y})$, an Inverse-Gamma as in (9); ..3.2 For $k=1,\dots,K$: ..3.2.1 For $l=1,\dots,p$: ..3.2.1.1 Sample ${\eta_{kl}^{2}}^{(c)}\sim f(\eta_{kl}^{2}|b_{kl}^{(c-1)},{\sigma^{2}}^{(c)})$, an Inverse-Normal as in (11); ..3.2.1.2 Set $\tau_{kl}^{2}=\frac{1}{\eta_{kl}^{2}}$; ..3.3 For $l=1,\dots,p$: ..3.3.1 Sample $\mu_{l}^{(c)}\sim f(\mu_{l}|\theta_{l}^{(c-1)})$, a continuous Bernoulli as in (12); ..3.3.2 Sample $Z_{l}^{(c)}\text{ with }\mathrm{P}(Z_{l}=1|\boldsymbol{b}^{(c-1)},\theta_{l}^{(c-1)},{\sigma^{2}}^{(c)},\boldsymbol{Z_{-[l]}}^{(*)},\boldsymbol{y})$ as in (8); ..3.3.3 Sample $\theta_{l}^{(c)}\sim f(\theta_{l}|\mu_{l}^{(c)},Z_{l}^{(c)})$, a beta as in (7); ..3.4 Sample $\boldsymbol{b}^{(c)}\sim f(\boldsymbol{b}|{\sigma^{2}}^{(c)},\boldsymbol{\tau^{2}}^{(c)},\boldsymbol{Z}^{(c)},\boldsymbol{y})$, a multivariate normal as in (6).
The term $\boldsymbol{Z_{-[l]}}^{(*)}$ in the panel 2.2 represents a mixed vector with components from the previous iteration that have not yet been updated and with components that have already been updated in the current iteration. As $l$ evolves in 3.3, the vector $\boldsymbol{Z_{-[l]}}^{(*)}$ is updated. Also, it is worth mentioning that the complete conditional distributions are presented in the panel in summary form, since the independent components are omitted from the notation.
@@
3 Numerical experiments
Numerical experiments were carried out with synthetic data in order to assess the main properties of the proposed model. In Section 3.1, we describe details about the synthetic data generation and the Gibbs sampler implementation. Section 3.2 presents the performance metrics used for model evaluation. Then, in Section 3.3, we present the results of our numerical experiments.
3.1 Synthetic data generation and Gibbs implementation
The synthetic data consists of $m=10$ (or $m=15$) curves generated as follows:

### v5

@@ 2 Proposed Bayesian model for variable selection in FOSR
The hyperparameter $\lambda$ controls both the mean and variance of $\tau_{kl}^{2}$, which in turn affects the prior variance of the coefficients.
@@ 2.1 Gibbs sampler
The panel 2.1 presents the standard procedure for implementing the Gibbs sampler in a clear and summarized way. Details on the derivation of each full conditional distribution are shown in Appendix A.
1. Define the hyperparameters. 2. Assign initial chain values to parameters. 3. For $\text{c}=2,\dots,\text{Nint}$: 3.1 Sample ${\sigma^{2}}^{(c)}\sim f(\sigma^{2}|\boldsymbol{b}^{(c-1)},\boldsymbol{\tau^{2}}^{(c-1)},\boldsymbol{Z}^{(c-1)},\boldsymbol{y})$, an inverse gamma as in (A.4). 3.2 For $k=1,\dots,K$: 3.2.1 For $l=1,\dots,p$: 3.2.1.1 Sample ${\eta_{kl}^{2}}^{(c)}\sim f(\eta_{kl}^{2}|b_{kl}^{(c-1)},{\sigma^{2}}^{(c)})$, an inverse normal as in (A.6). 3.2.1.2 Set $\tau_{kl}^{2}=1/\eta_{kl}^{2}$. 3.3 For $l=1,\dots,p$: 3.3.1 Sample $\mu_{l}^{(c)}\sim f(\mu_{l}|\theta_{l}^{(c-1)})$, a continuous Bernoulli as in (A.7). 3.3.2 Sample $Z_{l}^{(c)}$ with $\mathrm{P}(Z_{l}=1|\boldsymbol{b}^{(c-1)},\theta_{l}^{(c-1)},{\sigma^{2}}^{(c)},\boldsymbol{Z_{-[l]}}^{(*)},\boldsymbol{y})$ as in (A.3). 3.3.3 Sample $\theta_{l}^{(c)}\sim f(\theta_{l}|\mu_{l}^{(c)},Z_{l}^{(c)})$, a beta as in (A.2). 3.4 Sample $\boldsymbol{b}^{(c)}\sim f(\boldsymbol{b}|{\sigma^{2}}^{(c)},\boldsymbol{\tau^{2}}^{(c)},\boldsymbol{Z}^{(c)},\boldsymbol{y})$, a multivariate normal as in (A.1).
The term $\boldsymbol{Z_{-[l]}}^{(*)}$ in the panel 2.1 represents a mixed vector with components from the previous iteration that have not yet been updated and with components that have already been updated in the current iteration. As $l$ evolves in 3.3, the vector $\boldsymbol{Z_{-[l]}}^{(*)}$ is updated. Also, it is worth mentioning that the complete conditional distributions are presented in the panel in summary form, since the independent components are omitted from the notation. The definition of the continuous Bernoulli distribution can be found in (Loaiza-Ganem and Cunningham, 2019).
@@
3 Simulation study
Numerical experiments were carried out with simulated data to assess the main properties of the proposed model. To ensure an in-depth simulation study while avoiding an expensive computational cost, we deliberately opt to generate each simulated dataset considering $m=10$ curves and $n_{i}=n=25$ evaluation points per curve. This setup allows us to thoroughly explore the properties and potentialities of the proposed model. We further apply the proposed model using a larger dataset in Section 4. In addition, we simulated data considering two dispersion levels for the data ($\sigma=0.2$ and $\sigma=20$). We fitted each simulated dataset using our proposed Bayesian approach, considering $K=10$ B-spline basis functions for the expansion of the functional coefficients. We also investigated the sensitivity of our approach to the choice of $\mu$. Therefore, ten model fit configurations were tested under each dispersion scenario, nine of which considered $\mu$ as a hyperparameter and one that considered it as a parameter. For all model configurations, regardless of the dispersion scenario, a total of one hundred simulated datasets were used.
Although multiple values of the regularization parameter $\lambda$ could be tested, as is typically done for competing methods, we fixed $\lambda=\sqrt{2}$ (yielding a prior mean of one) due to the substantial computational cost of fitting the proposed model across one hundred replications and multiple configurations. Still, even with this fixed value, our proposed model showed excellent performance, with great fit and superior covariate selection accuracy.
With regard to the other hyperparameters, we set $\delta_{1}=\delta_{2}=0$. Thus, when using a degenerated prior to $\sigma^{2}$ with $\delta_{1}=0$ and $\delta_{2}=0$, there is an equivalence with the use of a non-informative improper prior $\frac{1}{\sigma^{2}}$. In situations where $\boldsymbol{\mu}$ is considered as a parameter, $\psi=0.6$ was defined.
Subsection 3.1 describes details about the simulated data generation and the Gibbs sampler implementation. Subsection 3.2 presents the performance metrics used for model evaluation. Then, in Subsection 3.3, we present the results of our numerical experiments.
3.1 Simulated data generation and Gibbs implementation
Each simulated dataset consists of $m=10$ curves generated as follows: